OpenStudy (abbycross167):

Can someone please show me how to do this? I really need help... I will give a medal Graph the following equations using the point-slope method or intercepts method. (1 pt) Show the steps of the procedure as outlined in the lesson as you graph the lines. (1 pt) Draw your y-intercept in yellow, your other points in red and your line in blue. y = (1/3)x – 1 Ok so I have done one like this before but the 1/3 confuses me... Please help me

OpenStudy (abbycross167):

@baru can you help me? the 1/3 is confusing

OpenStudy (dakid88):

y=-1/3x+1 Let x=0 and solve for y y=-1/3(0)+1 y=0+1 y=1 Plot (0,1) Let x=3 and solve for y y=-1/3(3)+1 y=-1+1 y=0 Plot (3,0) Connect the points and you have

OpenStudy (abbycross167):

okay Thank you very much!